The average income, I, in dollars, of a lawyer with an age of x years is modeled with the following function: I = -425x^2 + 45,500x – 650,000. according to this model, what is the predicted average annual income of 40-year old lawyers


1
Expert's answer
2020-08-10T18:11:48-0400

I= -425x2 + 45,500x - 650,000.

where x is the number of years.

therefore when x = 40

Then;

I= -425(40)2 + 45,500(40) - 650,000

I= -425(1600) + 45,500(40) - 650,000

I= -680,000 + 1,820,000 - 650,000

I= $ 490,000

The average annual income of a 40 year old is $ 490,000